# Population demographics of white dwarf binaries with intermediate separations: Gaia constraints on post-AGB mass transfer

### v1

@@ A.3 Discrepancy in the color excess distribution
As seen in Figure 5, the distribution of color excess is significantly more peaked around zero for our model compared to the observed sample. A potential cause is that the measured metallicities either (a) do not reflect the true values or (b) are drawn from a broader distribution than assumed in our Galaxia simulations. Here, we test this possibility by randomly drawing metallicities and their uncertainties from the observed non-class I sample in calculating $(B-I)_{\rm expected}$. Similarly to our fiducial model, we once again select 20% of the systems to have solar metallicity with an error of 0.25 dex. In Figure 18, we show the color excess distributions of the non-class I and NCE samples obtained after this change has been made. The mismatch introduced between the true and “measured” metallicities significantly broadens the color excess distribution compared to that in Figure 5 so that it is in better agreement with the observed sample. The broadening is also seen in panel D of Figure 17, which shows a 2D histogram of color excess and primary mass. We conclude that biases in the inferred metallicities for triples could plausibly explain both the broadness of the observed color excess distribution and the observed trend in color excess with metallicity. This possibility is also appealing because the metallicity distribution of the observed non-class I sample is significantly broader, and centered on more metal-poor values, than the metallicity distribution of all stars in the solar neighborhood (e.g. Hayden et al., 2015). However, we emphasize that this has minimal effect on the distributions of key stellar and orbital parameters in the final WD+MS binary sample plotted in Figure 7, because the triples with potentially biased metallicities are ultimately excluded from the final WD+MS binary sample.
@@
Appendix B Varying the $\beta$ function
As described in Section 5.1, our fiducial model assumes that systems that interact with an AGB donor evolve such that the final separation is a fraction $\beta$ of the initial separation which is a function of the initial mass ratio (Figure 10). This is chosen to be sigmoid function with the general functional form:
$\beta\left(q_{\rm init}\right)=\frac{A}{1+e^{-k(q_{\rm init}-q_{0})}}$ (B1)
where $A$ determines the asymptotic value to which the function maximizes, $k$ controls the steepness, and $q_{0}$ defines the critical mass ratio about which beta increases and the orbit remains wide. In our fiducial model, we set $A=0.63$, $k=60$, and $q_{0}=0.38$ (Figure 10). Here, we show the effects of varying each of these three parameters on the final distributions of the orbital period and component masses. On the leftmost panel of Figure 19, we plot the three additional $\beta$ functions tested. To the right, we show the parameter distributions of the final NCE samples obtained using these models. Once again, in order to more easily compare the shapes of the distributions, in these plots, the total numbers of sources in all samples have been adjusted to be equal.
We see that increasing the value of $A$ (i.e. reducing the amount of orbital shrinkage) proportionately increases the number of systems at shorter periods. This is because more systems in initially smaller orbits stay wide enough to receive orbital solutions. The total number of systems that enter the final sample is comparable to the fiducial model, and thus to the observed sample. The component mass distributions are also minimally changed from the fiducial model.
Reducing $k$, which makes the rise in $\beta$ shallower, leads to an orbital period distribution more steeply increasing towards shorter periods. As systems across a wider range of initial mass ratios remain in wide orbits and receive orbital solutions, the distributions of the component masses are slightly broader. This also leads to too many systems entering the final sample, with roughly three times more WD+MS binaries in the final sample than are in the observed sample.
Lastly, pushing $q_{0}$ to larger values primarily affects the component masses. As expected, the peak of the primary mass distribution occurs at a slightly larger value compared to the fiducial model and the deficit of high mass WDs is even more pronounced. As discussed in Section 6, this significantly reduces the total number of systems in the final sample, in this case to about half of that in our fiducial model, because systems must have higher initial mass ratios to end up in wide orbits.
Figure 19: The leftmost panel plots the $\beta$ functions tested. Each of the three additional models have one free parameter modified from our fiducial model (Section 5; Figure 10). The three panels to the right are analogous to those in Figure 15, but for the resulting final NCE samples obtained from using the modified $\beta$ functions. Note that in these plots, the total numbers have been adjusted to be the same across the models. Compared to our fiducial model, the modified versions perform worse in reproducing the observed period distribution and/or lead to too many or few systems entering the final sample.

### v3

@@ A.3 Discrepancy in the color excess distribution
As seen in Figure 5, the distribution of color excess is significantly more peaked around zero for our model compared to the observed sample. A potential cause is that the measured metallicities either (a) do not reflect the true values or (b) are drawn from a broader distribution than assumed in our Galaxia simulations. Here, we test this possibility by randomly drawing metallicities and their uncertainties from the observed non-class I sample in calculating $(B-I)_{\rm expected}$. Similarly to our empirical model, we once again select 20% of the systems to have solar metallicity with an error of 0.25 dex. In Figure 19, we show the color excess distributions of the non-class I and NCE samples obtained after this change has been made. The mismatch introduced between the true and “measured” metallicities significantly broadens the color excess distribution compared to that in Figure 5 so that it is in better agreement with the observed sample. The broadening is also seen in panel D of Figure 18, which shows a 2D histogram of color excess and primary mass. We conclude that biases in the inferred metallicities for triples could plausibly explain both the broadness of the observed color excess distribution and the observed trend in color excess with metallicity. This possibility is also appealing because the metallicity distribution of the observed non-class I sample is significantly broader, and centered on more metal-poor values, than the metallicity distribution of all stars in the solar neighborhood (e.g. Hayden et al., 2015). However, we emphasize that this has minimal effect on the distributions of key stellar and orbital parameters in the final WD+MS binary sample plotted in Figure 7, because the triples with potentially biased metallicities are ultimately excluded from the final WD+MS binary sample.
@@
A.4 WD IFMR
Figure 20: Left: Core mass as a function of initial mass and radius on the AGB. The markers are the data points from MIST tracks used in the interpolation. Right: The resulting astrometric sample of WD+MS binary candidates on the $P_{\rm orb}-M_{\rm 2,dark}$ from using the IFMR derived from interpolating the core masses in the left panel. A detailed description of the plot can be found in the text under Figure 8.
We also test the effect of using a different WD IFMR which takes into account the interruption of the progenitor’s evolution due to the onset of binary interaction. This may be particularly important if the system undergoes CEE and the envelope is quickly ejected. On average, this should lead to lower WD masses compared to those predicted from single star evolution and possibly ameliorate the discrepancy between the observed and simulated WD mass distributions (Section 4). We estimate WD mass as the he_core_mass from the MIST evolutionary tracks at a given giant radius and initial mass (Left panel of Figure 20; note that these cores are primarily carbon/oxygen by mass – the MESA label only indicates that the outer helium layer is included). We used the same method for the interpolation to predict the primary mass at the onset of MT (Section 5.4.1). We see that the core mass can exceed $1.4\,M_{\odot}$ at the final stages of the AGB phase for progenitor masses $\sim 6-8\,M_{\odot}$. In this case, we return to using the Weidemann (2000) IFMR. This is reasonable as the core is defined in terms of the helium mass fraction and not all of this material ends up burning into carbon and oxygen. A very small fraction of systems are in this regime so the end results are insensitive to this choice.
For stars that interact on the AGB, this IFMR leads to WD masses that are $\sim 10-20\%$ lower than those predicted by the Weidemann (2000) relation for progenitor masses between $\sim 1.5$ and $4.5\,M_{\odot}$. This corresponds to WD masses between $\sim 0.55-0.85\,M_{\odot}$. Outside this range, our IFMR can lead to higher WD masses. For comparison, Shahaf (2025) used a different method on a spectroscopic sample of AU-scale WD + red giant binaries in open clusters and found that these WDs were $\sim 20\%$ less massive than isolated WDs, which they attributed to binary interaction (also see Ironi et al. 2025).
The final WD+MS binary sample resulting from using this modified IFMR with the stable MT model (Section 5.4.1) is plotted on the $P_{\rm orb}-M_{\rm 2,dark}$ space in the right panel of Figure 20. All other assumptions are the same as in our empirical model (Section 5.4.4). We see that while the orbital period is not dramatically affected, the WD mass distribution now peaks at $\sim 0.55\,M_{\odot}$, below that of the observed sample. Furthermore, there remains too many WDs above $\sim 0.8\,M_{\odot}$.